OpenStudy (anonymous):

please someone help me with this. You plan to construct a brick wall measuring 15 ft. in length, 6 ft. high, and 1 ft. in width. If a brick measures 6 in. x 4 in. x 2 in., how many bricks will you need? (Remember to round up.) Total volume (to the nearest cubic inch) = __cubic inches. The number of bricks = __

OpenStudy (anonymous):

can u give the dimensions in metric system?

OpenStudy (anonymous):

12in=1 foot

OpenStudy (anonymous):

volume of wall = 15 *6*1 cubic feet volume of 1 brick =4*2*1 volume of "N " of bricks = volume of wall thr4 4*2*1 N =15*6*1 n=11.25 bricks = 11 bricks